OpenStudy (anonymous):

What is the inverse of f(x)=1/4x-2?

OpenStudy (anonymous):

\[y = \frac{1}{4}x - 2\] switch x's and y's \[x = \frac{1}{4}y - 2\] solve for y \[y = 4(x+2)\] replace y with \[f^{-1}(x)\] So the answer is \[f^{-1}(x) = 4(x+2)\]